40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 21 23:59:12 2018, shruti, Update, PEM, Seismometer temp control 
    Reply  Mon Jun 25 22:27:57 2018, shruti, Update, PEM, Seismometer temp control - heater circuit 
    Reply  Sun Jul 29 20:03:02 2018, rana, Update, PEM, Seismometer temp control Screen_Shot_2018-07-29_at_8.00.27_PM.png
Message ID: 14001     Entry time: Thu Jun 21 23:59:12 2018     Reply to this: 14016   14113
Author: shruti 
Type: Update 
Category: PEM 
Subject: Seismometer temp control 

We (Rana and I) are re-assembling the temperature controls on the seismometer to attempt PID control and then improve it using reinforcement learning.

We tried to re-assemble the connections for the heater and in-loop temperature sensor on the can that covers the seismometer.

We fixed (soldered) two of the connections from the heater circuit to the heater, but did not manage to get the PID working as one of the wires attached to the MOSFET had come off. Re-soldering the wire would be attempted tomorrow.

Equipment for undertaking all this is still left at the X-end of the interferometer and will be cleared soon.

ELOG V3.1.3-